### v2

# Low-degree mod 2 cohomology of classifying spaces of $G_{2}$-gauge groups

Abstract.
Let $G$ be a simply connected compact simple Lie group and let $\mathcal{G}_{k}$ denote the gauge group of a principal GG–bundle over $S^{4}$ with second Chern class $k\in\pi_{4}(BG)\cong\mathbb{Z}$. For $G=G_{2}$, the pp–local homotopy types of the gauge groups have been completely classified by Kishimoto–Theriault–Tsutaya and Kameko in terms of the order of the fundamental Samelson product $\langle i_{3},1\rangle\in[\Sigma^{3}G_{2},G_{2}]$.
In this paper, we begin a complementary study of the mod $2$ cohomology of the classifying spaces $B\mathcal{G}_{k}(G_{2})$. Our goal is to understand the structure of $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ as an unstable module over the mod $2$ Steenrod algebra in a low range of degrees. Using the evaluation fibration
$\Omega_{0}^{3}G_{2}\longrightarrow B\mathcal{G}_{k}\xrightarrow{\;\mathrm{ev}\;}BG_{2}$
together with Serre and Eilenberg–Moore spectral sequences, we study the Serre spectral sequence
$H^{s}(BG_{2};H^{t}(\Omega^{3}_{0}G_{2}))\Longrightarrow H^{s+t}(B\mathcal{G}_{k})$
in total degree $\leq 10$. A careful analysis of the homotopy groups of $G_{2}$ shows that
$H^{j}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})=0\quad\text{for }1\leq j\leq 4,\qquad H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})\neq 0,$
so the first positive-degree generator of the fibre cohomology occurs in degree $5$. As a consequence, there is a distinguished class
$u_{5}\in H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$
whose only possible Serre differential in total degree $\leq 10$ is a $d_{6}$–differential
$d_{6}(u_{5})=\varepsilon(k)\,x_{6}$
from $u_{5}$ to the degree-$6$ generator $x_{6}\in H^{6}(BG_{2};\mathbb{F}_{2})$, for a scalar $\varepsilon(k)\in\mathbb{F}_{2}$ encoding the low-degree effect of the bundle class. In addition, 22–locally we prove that $\varepsilon(k)$ is 44–periodic in $k$ (i.e. it depends only on $k\bmod 4$) and that $\varepsilon(k)=0$ for all $k\equiv 0\pmod{4}$.
Key words and phrases:
Gauge groups, Classifying spaces, Steenrod algebra, Serre spectral sequence
2020 Mathematics Subject Classification:
Primary 55R35; Secondary 55S10, 55P35, 55T10, 55T20
1. Introduction
1.1. Gauge groups and their homotopy types
Let $G$ be a compact, connected Lie group, let $X$ be a pointed CW complex, and let $P\to X$ be a principal GG–bundle. The gauge group $\mathcal{G}(P)$ is the topological group of GG–equivariant bundle automorphisms of $P$ that cover the identity on $X$. When $X$ is finite, Crabb–Sutherland showed that there are only finitely many homotopy types among the gauge groups $\mathcal{G}(P)$ as $P$ ranges over all principal GG–bundles over $X$ [4]. In the fundamental case $X=S^{4}$ and $G$ simply connected and simple, isomorphism classes of GG–bundles are classified by $\pi_{4}(BG)\cong\mathbb{Z}$, and we write $\mathcal{G}_{k}(G)$ for the gauge group of the bundle classified by $k\in\mathbb{Z}$.
The problem of determining the homotopy types of these gauge groups for specific Lie groups $G$ has received sustained attention; see for example [9, 12, 8]. In many cases the classification is governed by the order of a fundamental Samelson product in $G$.
The case $G=G_{2}$ is particularly interesting because of the exceptional nature of $G_{2}$. Kishimoto, Theriault and Tsutaya proved that the classification for $G_{2}$–gauge groups reduces to determining the order of the Samelson product
$\langle i_{3},1\rangle\in[\Sigma^{3}G_{2},G_{2}],$
where $i_{3}\colon S^{3}\to G_{2}$ is the inclusion of the bottom cell and $1$ denotes the identity map [8].
Very recently, Kameko [7] resolved this issue at the prime $2$ and completed the pp–local classification of $G_{2}$–gauge groups. He proved that the order of $\langle i_{3},1\rangle$ is exactly $84$ [7, Theorem 1.1]. Consequently, the pp–local homotopy type of $\mathcal{G}_{k}(G_{2})$ depends only on the greatest common divisor $(k,84)$. Thus the homotopy types of the gauge groups are completely understood up to pp–localization. It is natural to ask how far this information is reflected in the (co)homology of their classifying spaces.
1.2. Cohomology of $B\mathcal{G}_{k}$ and the Steenrod algebra
Gottlieb identified the classifying space of $\mathcal{G}_{k}$ with a component of a mapping space [5]: specifically, $B\mathcal{G}_{k}$ is homotopy equivalent to $\mathrm{Map}_{k}(S^{4},BG)$, the connected component of the mapping space $\mathrm{Map}(S^{4},BG)$ containing the classifying map $k\colon S^{4}\to BG$. This identification yields a fundamental evaluation fibration
$\Omega^{3}_{0}G\;\longrightarrow\;B\mathcal{G}_{k}\xrightarrow{\;\mathrm{ev}\;}BG.$
For $G=G_{2}$, this expresses $B\mathcal{G}_{k}$ as a twisted mapping space over the base $BG_{2}$.
The mod $2$ cohomology of $BG_{2}$ is classically known: Borel showed that
$H^{*}(BG_{2};\mathbb{F}_{2})\cong\mathbb{F}_{2}[x_{4},x_{6},x_{7}]$
is a polynomial algebra on generators in degrees $4,6,7$ [2]. The action of the Steenrod algebra $\mathcal{A}$ on $H^{*}(BG_{2})$ can be described explicitly via Wu formulas, and from the viewpoint of the hit problem one is interested in minimal generating sets for the unstable $\mathcal{A}$–module $\mathbb{F}_{2}[x_{4},x_{6},x_{7}]$.
By contrast, there seems to be no published description of the full mod $2$ cohomology ring $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ for any nontrivial $k$, nor of its structure as an unstable $\mathcal{A}$–module. Our aim here is to give a precise structural description of $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ in low degrees, in terms of the evaluation fibration and the first nontrivial cohomology of the triple loop space $\Omega^{3}_{0}G_{2}$.
1.3. Main result: a low-degree description
We work 22–locally throughout and write $H^{*}(-)$ for $H^{*}(-;\mathbb{F}_{2})$. Consider the Serre spectral sequence of the evaluation fibration
(1.1) $E_{2}^{s,t}(k)\;\cong\;H^{s}(BG_{2};H^{t}(\Omega^{3}_{0}G_{2}))\;\Longrightarrow\;H^{s+t}(B\mathcal{G}_{k}).$
The $E_{2}$–term is a module over $H^{*}(BG_{2})$ and an unstable module over the Steenrod algebra. The twisting of this spectral sequence depends on $k$ through the connecting map
$\partial_{k}\colon G_{2}\longrightarrow\Omega^{3}_{0}G_{2},$
which is the triple adjoint of the Samelson product $\langle k\cdot i_{3},1\rangle$ by work of Kishimoto–Theriault–Tsutaya [8]. Kameko’s calculation of the 22–primary order of $\langle i_{3},1\rangle$ implies that, after localization at $2$, the map $\partial_{k}$ depends only on $k\bmod 4$.
The cohomology of the fibre $\Omega^{3}_{0}G$ can in principle be computed by iterated Eilenberg–Moore spectral sequences applied to the path-loop fibrations of $BG$ and $G$. Choi and Yoon conjectured that, for any simply connected finite HH–space $X$, the Eilenberg–Moore spectral sequences converging to the mod $p$ homology of $\Omega^{2}X$ and $\Omega^{3}X$ collapse at $E_{2}$ [3]. This conjecture was subsequently proved by Lin [10]; in particular, for any compact simple Lie group $G$ and any prime $p$, the Eilenberg–Moore spectral sequences computing $H_{*}(\Omega^{2}G;\mathbb{F}_{p})$ and $H_{*}(\Omega^{3}G;\mathbb{F}_{p})$ collapse at $E_{2}$. Combined with the classical description of $H^{*}(BG_{2})$, this implies that $H^{*}(\Omega^{3}_{0}G_{2})$ is of finite type in each degree and that its generators occur in bounded degrees. For our purposes, however, we only need to know the very first nontrivial cohomology group of $\Omega^{3}_{0}G_{2}$.
Let
$M\;=\;H^{*}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$
and write $M^{\leq N}=\bigoplus_{j\leq N}M^{j}$ for its truncation in degrees $\leq N$. Our main theorem describes the Serre spectral sequence (1.1) in total degree $\leq 10$ and singles out a distinguished scalar $\varepsilon(k)\in\mathbb{F}_{2}$ which already captures the effect of $k$ on the first nontrivial fibre class.
Theorem 1.1 (Low-degree structure of $H^{*}(B\mathcal{G}_{k})$).
Work 22–locally and fix an integer $k$. Consider the Serre spectral sequence (1.1) for the fibration $\Omega^{3}_{0}G_{2}\to B\mathcal{G}_{k}\to BG_{2}$, and let $M=H^{*}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$.
  1. (i)
    In total degree $j=s+t\leq 10$ the $E_{2}$–term is naturally isomorphic, as a bigraded $H^{*}(BG_{2})$–module, to
    $E_{2}^{*,*}(k)^{\leq 10}\;\cong\;H^{*}(BG_{2})\otimes M^{\leq 10},$
    with $E_{2}^{s,t}=H^{s}(BG_{2})\otimes M^{t}$.
  2. (ii)
    Using the low-degree homotopy groups of $G_{2}$ and the Hurewicz theorem, one has
    $M^{j}=H^{j}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})=0\quad\text{for }1\leq j\leq 4,\qquad M^{5}\neq 0.$
    In particular, there exists a nonzero class
    $u_{5}\in M^{5}\cong E_{2}^{0,5}(k)$
    and $u_{5}$ is the first positive-degree generator of $H^{*}(\Omega^{3}_{0}G_{2})$.
  3. (iii)
    In total degree $j\leq 10$ the differentials
    $d_{r}\colon E_{r}^{0,t}(k)\longrightarrow E_{r}^{r,t-r+1}(k)$
    with $2\leq r\leq 5$ vanish on all classes in bidegree $(0,t)$, and the only page on which a differential with source in $E_{r}^{0,5}(k)$ can be nonzero is $r=6$. After choosing a generator $x_{6}\in H^{6}(BG_{2})\cong E_{2}^{6,0}(k)$ and $u_{5}\in M^{5}\cong E_{2}^{0,5}(k)$ there is a uniquely determined scalar $\varepsilon(k)\in\mathbb{F}_{2}$ such that
    $d_{6}(u_{5})=\varepsilon(k)\,x_{6}$
    in $E_{6}^{6,0}(k)$.
  4. (iv)
    Let $U\subset E_{2}^{*,*}(k)$ denote the $H^{*}(BG_{2})$–submodule generated by $u_{5}$. In total degree $j\leq 10$ all differentials $d_{r}$ with $2\leq r\leq 5$ vanish on $U$, and every differential
    $d_{6}\colon E_{6}^{s,5}(k)\longrightarrow E_{6}^{s+6,0}(k)$
    with $s+5\leq 10$ satisfies
    $d_{6}(x\cdot u_{5})=x\cdot d_{6}(u_{5})\qquad\text{for }x\in H^{s}(BG_{2}).$
    In particular, within $U$ the entire effect of the Serre differentials in total degree $\leq 10$ is determined by the single scalar $\varepsilon(k)\in\mathbb{F}_{2}$.
  5. (v)
    For each $j\leq 10$ the spectral sequence induces a finite filtration of $H^{j}(B\mathcal{G}_{k})$ whose associated graded contains, as a direct summand, the image of $U$ in $E_{\infty}^{*,*}(k)^{j}$. This summand is explicitly computable from $H^{*}(BG_{2})\otimes M^{\leq 10}$ together with $\varepsilon(k)$, while the remaining extension term depends only on $M^{\leq j}$ and the multiplicative structure of the Serre spectral sequence.
In particular, in total degree $j\leq 10$ the dependence of $H^{j}(B\mathcal{G}_{k};\mathbb{F}_{2})$ on the bundle class $k$ that is visible through the first nontrivial fibre class $u_{5}$ is completely encoded in the single Serre differential $d_{6}(u_{5})=\varepsilon(k)x_{6}$.
Remark 1.2.
Theorem 1.1 should be viewed as a structural reduction. It does not claim a closed formula for $\varepsilon(k)$, nor does it assert that $\varepsilon(k)$ is nontrivial. Kameko’s result shows that the 22–local homotopy type of $\mathcal{G}_{k}(G_{2})$ depends on $k\bmod 4$ through the 22–primary order of the Samelson product $\langle i_{3},1\rangle$ [8, 7]. It is therefore natural to expect that $\varepsilon(k)$ is governed by the same 22–primary data.
In fact, we prove (Lemma 4.5) that, 22–locally, the scalar $\varepsilon(k)$ depends only on the residue class of $k$ modulo $4$, and moreover $\varepsilon(k)=0$ whenever $4\mid k$ (Corollary 4.6). Thus the remaining low-degree problem reduces to determining the two values $\varepsilon(1)$ and $\varepsilon(2)$.
Determining these values explicitly would require a more detailed understanding of $H^{*}(\Omega^{3}_{0}G_{2})$ and of the induced map of the connecting map $\partial_{k}$ on mod $2$ cohomology, and we leave this as a natural problem for future work.
The paper is organised as follows. In Section 2 we review the necessary background on gauge groups, Samelson products, the mod $2$ cohomology of $BG_{2}$ and the Steenrod algebra. In Section 3 we describe the evaluation fibration and its Serre spectral sequence, and we recall the relationship between the connecting map and the Samelson product. Finally, Section 4 contains the proof of Theorem 1.1, including a detailed analysis of the bidegrees for $j\leq 10$ and the vanishing of $H^{j}(\Omega^{3}_{0}G_{2})$ for $1\leq j\leq 4$.
2. Preliminaries
2.1. Gauge groups and mapping spaces
Let $G$ be a compact, connected Lie group and $P\to S^{4}$ a principal GG–bundle. The associated gauge group $\mathcal{G}(P)$ is the group of GG–equivariant automorphisms of $P$ covering the identity on $S^{4}$, with the compact-open topology. When $G$ is simply connected and simple, its classifying space $BG$ is 33–connected and $\pi_{4}(BG)\cong\mathbb{Z}$; the integer $k\in\mathbb{Z}$ corresponding to $[P]$ plays the role of the second Chern class.
Following the notation of [8, 7], we write $\mathcal{G}_{k}=\mathcal{G}_{k}(G_{2})$ for the gauge group of the principal $G_{2}$–bundle classified by $k\in\pi_{4}(BG_{2})\cong\mathbb{Z}$. By a classical result of Gottlieb [5], the classifying space $B\mathcal{G}_{k}$ is homotopy equivalent to $\mathrm{Map}_{k}(S^{4},BG_{2})$, the connected component of the mapping space $\mathrm{Map}(S^{4},BG_{2})$ containing maps of degree $k$.
Under this identification, the evaluation map $\mathrm{ev}\colon\mathrm{Map}(S^{4},BG_{2})\to BG_{2}$ restricts to a fibration
$\mathrm{Map}_{*}(S^{4},BG_{2})_{k}\longrightarrow B\mathcal{G}_{k}\xrightarrow{\;\mathrm{ev}\;}BG_{2},$
where the fibre $\mathrm{Map}_{*}(S^{4},BG_{2})_{k}$ is homotopy equivalent to the triple loop space $\Omega^{3}_{0}G_{2}$.
When $G$ is simply connected, the based mapping space $\mathrm{Map}_{*}(S^{4},BG)_{k}$ is homotopy equivalent to the triple loop space $\Omega^{3}_{0}G$ for every $k$, and via Gottlieb’s equivalence we obtain the fibration
$\Omega^{3}_{0}G\;\longrightarrow\;B\mathcal{G}_{k}(G)\xrightarrow{\;\mathrm{ev}\;}BG.$
We henceforth specialise to $G=G_{2}$.
2.2. Samelson products and the connecting map
Let $G$ be an HH–space. The Samelson product
$\langle-,-\rangle\colon\pi_{m}(G)\otimes\pi_{n}(G)\longrightarrow\pi_{m+n}(G)$
is defined by the composite
$S^{m}\wedge S^{n}\xrightarrow{f\wedge g}G\wedge G\xrightarrow{[-,-]}G,$
where $f,g$ represent the chosen homotopy classes and [,][-,-] is the commutator in $G$. For a compact simple Lie group $G$ the order of certain Samelson products controls the classification of gauge groups; see for example [9, 12, 8] for concrete instances of this phenomenon.
In the present context the relevant Samelson product is
$\langle i_{3},1\rangle\in[\Sigma^{3}G,G],$
where $i_{3}\colon S^{3}\to G$ is the inclusion of the bottom cell and $1$ is the identity. Kishimoto, Theriault and Tsutaya showed that the connecting map
$\partial_{k}\colon G\longrightarrow\Omega^{3}_{0}G$
in the homotopy fibration
$\mathcal{G}_{k}(G)\;\longrightarrow\;G\xrightarrow{\;\partial_{k}\;}\Omega^{3}_{0}G\longrightarrow B\mathcal{G}_{k}(G)\xrightarrow{\;\mathrm{ev}\;}BG$
is the triple adjoint of the Samelson product $\langle k\cdot i_{3},1\rangle$ [8, Lemma 2.1]. In particular $\partial_{k}\simeq k\cdot\partial_{1}$, and the order of $\partial_{1}$ is equal to the order of $\langle i_{3},1\rangle$.
For $G=G_{2}$, the work of Kishimoto–Theriault–Tsutaya [8] and Kameko [7] implies that the 22–primary component of $\langle i_{3},1\rangle$ has order $4$. Thus, after localizing at $2$, the homotopy class of $\partial_{k}$ depends only on $k\bmod 4$. This 22–primary information is the only input from the classification of $G_{2}$–gauge groups that will enter our cohomological analysis.
3. The evaluation fibration and its spectral sequence
3.1. The Serre spectral sequence
Specialising the evaluation fibration to $G=G_{2}$, we obtain a fibration of simply connected spaces
$\Omega^{3}_{0}G_{2}\;\longrightarrow\;B\mathcal{G}_{k}\xrightarrow{\;\mathrm{ev}\;}BG_{2}$
and hence a mod $2$ cohomology Serre spectral sequence
$E_{2}^{s,t}(k)=H^{s}(BG_{2};H^{t}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2}))\;\Longrightarrow\;H^{s+t}(B\mathcal{G}_{k};\mathbb{F}_{2}).$
We abbreviate $M=H^{*}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$; it is a graded commutative algebra and an unstable $\mathcal{A}$–module.
Since $BG_{2}$ is simply connected and the action of $\pi_{1}(BG_{2})$ on $H^{*}(\Omega^{3}_{0}G_{2})$ is trivial, the local coefficient system on the fibre is trivial, and we have
$E_{2}^{s,t}(k)\cong H^{s}(BG_{2})\otimes M^{t}$
as bigraded vector spaces, with multiplicative structure induced from the tensor product.
The spectral sequence is natural with respect to maps of fibrations. In particular, the connecting map $\partial_{k}\colon G_{2}\to\Omega^{3}_{0}G_{2}$ in the associated homotopy fibration controls the kk–dependence of the differentials.
3.2. The triple loop space $\Omega^{3}_{0}G_{2}$
The homotopy groups of $G_{2}$ in low degrees were computed by Mimura and Toda [11]. For our purposes we only need the following facts:
$\pi_{3}(G_{2})\cong\mathbb{Z},\qquad\pi_{4}(G_{2})=0,\qquad\pi_{5}(G_{2})=0,\qquad\pi_{6}(G_{2})\text{ is $3$--torsion},\qquad\pi_{7}(G_{2})=0,$
and $\pi_{8}(G_{2})$ contains 22–torsion. The precise structure of $\pi_{6}(G_{2})$ and $\pi_{8}(G_{2})$ will not be needed.
We focus on the component $\Omega^{3}_{0}G_{2}$ containing the constant map.
Lemma 3.1.
The space $\Omega^{3}_{0}G_{2}$ is 22–connected. In particular
$\pi_{1}(\Omega^{3}_{0}G_{2})=\pi_{2}(\Omega^{3}_{0}G_{2})=0.$
Proof.
By definition, $\Omega^{3}_{0}G_{2}$ denotes the path component of the constant map in the triple loop space $\Omega^{3}G_{2}$, so $\pi_{i}(\Omega^{3}_{0}G_{2})\cong\pi_{i}(\Omega^{3}G_{2})$ for all $i\geq 1$ and $\pi_{0}(\Omega^{3}_{0}G_{2})=0$.
For any based space $X$ and integer $n\geq 0$, there is a natural isomorphism
$\pi_{n}(\Omega X)\;\cong\;\pi_{n+1}(X),$
obtained by identifying both sides with homotopy classes of based maps via the suspension loop adjunction
[…]
Iterating this identification three times gives, for every $i\geq 0$,
$\pi_{i}(\Omega^{3}G_{2})\;\cong\;\pi_{i+1}(\Omega^{2}G_{2})\;\cong\;\pi_{i+2}(\Omega G_{2})\;\cong\;\pi_{i+3}(G_{2}).$
Since $\Omega^{3}_{0}G_{2}$ is the identity component of $\Omega^{3}G_{2}$, we in particular obtain
$\pi_{i}(\Omega^{3}_{0}G_{2})\;\cong\;\pi_{i+3}(G_{2})\qquad\text{for }i\geq 1.$
The low-degree homotopy groups of $G_{2}$ were computed by Mimura–Toda [11]; in particular
$\pi_{4}(G_{2})=0,\qquad\pi_{5}(G_{2})=0.$
Substituting $i=1,2$ in the isomorphism above yields
$\pi_{1}(\Omega^{3}_{0}G_{2})\cong\pi_{4}(G_{2})=0,\qquad\pi_{2}(\Omega^{3}_{0}G_{2})\cong\pi_{5}(G_{2})=0.$
Thus $\Omega^{3}_{0}G_{2}$ is 22–connected, as claimed. ∎
We can now determine the first few homology and cohomology groups of $\Omega^{3}_{0}G_{2}$ using the Hurewicz theorem and universal coefficients.
Proposition 3.2.
Let $M=H^{*}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$. Then:
  1. (a)
    $M^{0}\cong\mathbb{F}_{2}$ and
    $M^{1}=M^{2}=M^{3}=M^{4}=0.$
  2. (b)
    The first positive-degree cohomology group $M^{5}$ is nonzero:
    $M^{5}=H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})\neq 0.$
    In particular, there exists a nonzero class $u_{5}\in M^{5}$.
Proof.
We work integrally and then pass to mod $2$ coefficients.
Step 1: Low-degree homology of $\Omega^{3}_{0}G_{2}$. By Lemma 3.1, the space $X=\Omega^{3}_{0}G_{2}$ is 22–connected. The Hurewicz theorem for (n1)(n-1)–connected spaces states that if $X$ is (n1)(n-1)–connected with $n\geq 2$, then $H_{i}(X;\mathbb{Z})=0$ for $1\leq i\leq n-1$, and the Hurewicz map
$h\colon\pi_{n}(X)\longrightarrow H_{n}(X;\mathbb{Z})$
is an isomorphism (see, for example, [6, Theorem 4.32]). In our case $n=3$, so $X$ is 22–connected and we obtain:
$H_{1}(X;\mathbb{Z})=H_{2}(X;\mathbb{Z})=0,\qquad h\colon\pi_{i}(X)\xrightarrow{\cong}H_{i}(X;\mathbb{Z})\quad\text{for }i=3,4,5.$
Using $\pi_{i}(X)\cong\pi_{i+3}(G_{2})$ for $i\geq 0$ and the low-degree homotopy groups of $G_{2}$ quoted above, we have
$H_{3}(X;\mathbb{Z})$ $\cong\pi_{3}(X)\cong\pi_{6}(G_{2})\quad\text{(a finite $3$--group)},$
$H_{4}(X;\mathbb{Z})$ $\cong\pi_{4}(X)\cong\pi_{7}(G_{2})=0,$
$H_{5}(X;\mathbb{Z})$ $\cong\pi_{5}(X)\cong\pi_{8}(G_{2}),$
and $\pi_{8}(G_{2})$ contains 22–torsion by the computations of Mimura and Toda [11]. Thus $H_{5}(X;\mathbb{Z})$ has nontrivial 22–primary torsion.
Step 2: Low-degree cohomology with $\mathbb{F}_{2}$–coefficients. We now pass to mod $2$ cohomology via the universal coefficient short exact sequence
$0\longrightarrow\operatorname{Ext}^{1}_{\mathbb{Z}}(H_{i-1}(X;\mathbb{Z}),\mathbb{F}_{2})\longrightarrow H^{i}(X;\mathbb{F}_{2})\longrightarrow\operatorname{Hom}_{\mathbb{Z}}(H_{i}(X;\mathbb{Z}),\mathbb{F}_{2})\longrightarrow 0.$
For $i=1$ we use $H_{0}(X;\mathbb{Z})\cong\mathbb{Z}$ and $H_{1}(X;\mathbb{Z})=0$ to obtain $H^{1}(X;\mathbb{F}_{2})=0$. For $i=2$ we use $H_{1}(X;\mathbb{Z})=H_{2}(X;\mathbb{Z})=0$ and obtain $H^{2}(X;\mathbb{F}_{2})=0$. For $i=3$ we have $H_{2}(X;\mathbb{Z})=0$, so
$H^{3}(X;\mathbb{F}_{2})\;\cong\;\operatorname{Hom}_{\mathbb{Z}}(H_{3}(X;\mathbb{Z}),\mathbb{F}_{2}).$
Since $H_{3}(X;\mathbb{Z})\cong\pi_{6}(G_{2})$ is a finite 33–group, and $3$ is invertible in $\mathbb{F}_{2}$, any homomorphism $H_{3}(X;\mathbb{Z})\to\mathbb{F}_{2}$ must be trivial, hence $H^{3}(X;\mathbb{F}_{2})=0$.
For $i=4$ we have $H_{3}(X;\mathbb{Z})$ a 33–group and $H_{4}(X;\mathbb{Z})=0$, so
$H^{4}(X;\mathbb{F}_{2})\;\cong\;\operatorname{Ext}^{1}_{\mathbb{Z}}(H_{3}(X;\mathbb{Z}),\mathbb{F}_{2}).$
For a cyclic group $\mathbb{Z}/n$ one has
$\operatorname{Ext}^{1}_{\mathbb{Z}}(\mathbb{Z}/n,\mathbb{F}_{2})\cong\mathbb{F}_{2}/n\mathbb{F}_{2},$
which is zero whenever $n$ is odd. Since $H_{3}(X;\mathbb{Z})$ is a 33–group, it follows that $\operatorname{Ext}^{1}_{\mathbb{Z}}(H_{3}(X;\mathbb{Z}),\mathbb{F}_{2})=0$, hence $H^{4}(X;\mathbb{F}_{2})=0$.
This proves $M^{1}=M^{2}=M^{3}=M^{4}=0$; the equality $M^{0}\cong\mathbb{F}_{2}$ follows from path-connectedness. This establishes (a).
For $i=5$ we use the fact that $H_{4}(X;\mathbb{Z})=0$, so
$H^{5}(X;\mathbb{F}_{2})\;\cong\;\operatorname{Hom}_{\mathbb{Z}}(H_{5}(X;\mathbb{Z}),\mathbb{F}_{2}).$
Since $H_{5}(X;\mathbb{Z})\cong\pi_{8}(G_{2})$ contains 22–torsion, there is a nonzero homomorphism $H_{5}(X;\mathbb{Z})\to\mathbb{F}_{2}$, and hence $H^{5}(X;\mathbb{F}_{2})\neq 0$. This proves (b). ∎
For the remainder of the paper we fix a nonzero element $u_{5}\in M^{5}=H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$.
4. Proof of the low-degree theorem
In this section we analyse the Serre spectral sequence (1.1) in total degree $j=s+t\leq 10$ and prove Theorem 1.1. We work throughout with mod $2$ coefficients and suppress them from the notation.
4.1. The $E_{2}$–page in low degrees
By the discussion in Section 3, the $E_{2}$–term has the form
$E_{2}^{s,t}(k)\cong H^{s}(BG_{2})\otimes M^{t},$
and $H^{*}(BG_{2})\cong\mathbb{F}_{2}[x_{4},x_{6},x_{7}]$ is generated by $x_{4},x_{6},x_{7}$ in degrees $4,6,7$. For $0<j\leq 10$ the only degrees $s>0$ in which $H^{s}(BG_{2})$ can be nonzero are
$s\in\{4,6,7,8,10\},$
corresponding to monomials $x_{4}$, $x_{6}$, $x_{7}$, $x_{4}^{2}$ and $x_{4}x_{6}$, respectively.
By Proposition 3.2, the fibre cohomology $M^{t}$ vanishes for $1\leq t\leq 4$. Consequently, in total degree $j=s+t\leq 10$ the only nonzero groups $E_{2}^{s,t}$ with $t>0$ occur when
$t\geq 5.$
In particular, the class $u_{5}\in M^{5}$ gives rise to a class
$u=1\otimes u_{5}\in E_{2}^{0,5}(k),$
and this is the first positive-degree generator of $M$.
4.2. Bidegree constraints on differentials
The differentials in the Serre spectral sequence have bidegree
$d_{r}\colon E_{r}^{s,t}\longrightarrow E_{r}^{s+r,t-r+1}.$
We now examine which values of $r$ can occur with source or target in total degree $j\leq 10$.
Let $(s,t)$ be a bidegree with $s+t\leq 10$ and $E_{r}^{s,t}\neq 0$. If $d_{r}$ is nonzero on $E_{r}^{s,t}$, then its target lies in bidegree (s+r,tr+1)(s+r,t-r+1) with total degree
$(s+r)+(t-r+1)=s+t+1\leq 11.$
Moreover, for $s>0$ we know that $H^{s}(BG_{2})$ can only be nonzero in degrees
$s\in\{4,6,7,8,10\},$
and by Proposition 3.2 the fibre cohomology satisfies
$M^{0}\cong\mathbb{F}_{2},\qquad M^{j}=0\ \text{for}\ 1\leq j\leq 4,\qquad M^{5}\neq 0.$
In this subsection we restrict attention to differentials whose source is the first nontrivial fibre class
$u_{5}\in M^{5}\cong E_{2}^{0,5}(k),$
and we determine on which page $u_{5}$ can possibly support a nontrivial differential in total degree $j\leq 10$.
Since $u_{5}$ lies in bidegree $(0,5)$, any differential with source $u_{5}$ has the form
$d_{r}(u_{5})\in E_{r}^{r,\,5-r+1}.$
We analyse the possible values of $r$.
  • If $r\geq 7$, then $5-r+1\leq-1$, so the target bidegree has negative fibre degree and $E_{r}^{r,5-r+1}=0$. Thus $d_{r}(u_{5})=0$ for all $r\geq 7$ in total degree $\leq 10$.
  • For $2\leq r\leq 5$ we have the following possibilities:
    • $r=2$: the target is $(2,4)$, but $H^{2}(BG_{2})=0$, so $E_{2}^{2,4}=0$ and hence $d_{2}(u_{5})=0$.
    • $r=3$: the target is $(3,3)$, and again $H^{3}(BG_{2})=0$, so $d_{3}(u_{5})=0$.
    • $r=4$: the target is $(4,2)$; here $H^{4}(BG_{2})\cong\mathbb{F}_{2}\{x_{4}\}$ is nonzero, but $M^{2}=0$ by Proposition 3.2, so $E_{2}^{4,2}\cong H^{4}(BG_{2})\otimes M^{2}=0$, and hence $d_{4}(u_{5})=0$.
    • $r=5$: the target is $(5,1)$, and $H^{5}(BG_{2})=0$, so $E_{2}^{5,1}=0$ and $d_{5}(u_{5})=0$.
    In particular, $d_{r}(u_{5})=0$ for all $2\leq r\leq 5$.
  • For $r=6$ the target is bidegree $(6,0)$, and
    $E_{2}^{6,0}\cong H^{6}(BG_{2})\otimes M^{0}\cong H^{6}(BG_{2})\cong\mathbb{F}_{2}\{x_{6}\}$
    is nonzero. There is therefore no a priori bidegree reason to force $d_{6}(u_{5})$ to vanish; this is the first page on which $u_{5}$ can possibly support a nontrivial differential in total degree $\leq 10$.
We summarise the discussion in the following form, which will be sufficient for our later arguments.
Lemma 4.1.
Let $u_{5}\in E_{2}^{0,5}(k)\cong M^{5}$ be a nonzero class representing the first positive-degree fibre cohomology. In total degree $j\leq 10$ all differentials $d_{r}(u_{5})$ vanish for $2\leq r\leq 5$ and for $r\geq 7$, and the only page on which $u_{5}$ can support a nontrivial differential is $r=6$. More precisely,
$d_{6}(u_{5})\in E_{6}^{6,0}(k)\cong H^{6}(BG_{2}).$
Fix once and for all a generator $x_{6}\in H^{6}(BG_{2})\cong\mathbb{F}_{2}$. Since $H^{6}(BG_{2})$ is one-dimensional over $\mathbb{F}_{2}$, there exists a uniquely determined scalar $\varepsilon(k)\in\mathbb{F}_{2}$ such that
$d_{6}(u_{5})=\varepsilon(k)\,x_{6}.$
Extending this analysis slightly, we obtain a convenient description of the possible differentials on the $H^{*}(BG_{2})$–submodule generated by $u_{5}$ in low total degrees.
Lemma 4.2.
In total degree $s+t\leq 10$ the following hold.
  1. (a)
    For any $r\geq 7$, every differential
    $d_{r}\colon E_{r}^{s,t}(k)\longrightarrow E_{r}^{s+r,t-r+1}(k)$
    has either source or target lying in total degree >10>10. In particular, no $d_{r}$ with $r\geq 7$ affects the groups $E_{r}^{s,t}(k)$ with $s+t\leq 10$.
  2. (b)
    If we restrict to the $H^{*}(BG_{2})$–submodule generated by $u_{5}\in E_{2}^{0,5}(k)$, then all differentials $d_{r}$ with $2\leq r\leq 5$ vanish in total degree $s+t\leq 10$, and every differential
    $d_{6}\colon E_{6}^{s,t}(k)\longrightarrow E_{6}^{s+6,t-5}(k)$
    in this range is determined multiplicatively by the value of $d_{6}$ on $u_{5}$. Concretely, for $s+t\leq 10$ and any $x\in H^{s}(BG_{2})$ we have
    $d_{6}(x\cdot u_{5})=x\cdot d_{6}(u_{5})\in E_{6}^{s+6,0}(k),$
    and there are no other possible nonzero differentials on this submodule. In particular, up to multiplication by elements of $H^{*}(BG_{2})$, the only potentially nonzero differential in the range $s+t\leq 10$ with source in this submodule is
    $d_{6}\colon E_{6}^{0,5}(k)\longrightarrow E_{6}^{6,0}(k).$
Proof.
We examine all possible differentials
$d_{r}\colon E_{r}^{s,t}(k)\longrightarrow E_{r}^{s+r,t-r+1}(k)$
with $s+t\leq 10$.
For part (a), suppose $r\geq 7$ and $E_{r}^{s,t}\neq 0$ with $s+t\leq 10$. Then the target bidegree (s+r,tr+1)(s+r,t-r+1) has total degree
$(s+r)+(t-r+1)=s+t+1\leq 11.$
If $s+t\leq 10$ and $r\geq 7$, a short check shows that as soon as $s+r\geq 11$ or tr+1<0t-r+1<0 the target lies either in total degree >10>10 or in negative fibre degree. In each case the target group is zero in our range of interest, so no such $d_{r}$ can hit or originate from classes with $s+t\leq 10$. This proves (a).
For part (b) we restrict to the $H^{*}(BG_{2})$–submodule generated by $u_{5}$, i.e. to classes of the form
$x\cdot u_{5}\in E_{2}^{s,5}(k),\qquad x\in H^{s}(BG_{2}),$
with $s+5\leq 10$. Since $H^{s}(BG_{2})$ is nonzero only for $s\in\{0,4,6,7,8,10,\dots\}$, the only possibilities with $s+5\leq 10$ are $s=0$ and $s=4$, giving bidegrees $(0,5)$ and $(4,5)$.
We first analyse the differentials on $u_{5}$ itself. By Lemma 4.1 we know that $d_{r}(u_{5})=0$ for $2\leq r\leq 5$ and for $r\geq 7$, and that the only page on which a nontrivial differential can occur is $r=6$, with
$d_{6}(u_{5})\in E_{6}^{6,0}(k)\cong H^{6}(BG_{2}).$
Now let $x\in H^{s}(BG_{2})$ with $s+5\leq 10$ and consider the class $x\cdot u_{5}\in E_{2}^{s,5}(k)$.
(i) The case $2\leq r\leq 5$. Any differential with source $x\cdot u_{5}$ has target in bidegree (s+r,5r+1)(s+r,5-r+1). For $2\leq r\leq 5$ we have $5-r+1\in\{4,3,2,1\}$, so by Proposition 3.2 the fibre term $M^{5-r+1}$ vanishes and hence $E_{2}^{s+r,5-r+1}(k)=0$. Thus $d_{r}(x\cdot u_{5})=0$ for all $2\leq r\leq 5$.
(ii) The case $r=6$. Here the target of $d_{6}$ is
$d_{6}\colon E_{6}^{s,5}(k)\longrightarrow E_{6}^{s+6,0}(k)\cong H^{s+6}(BG_{2}).$
The Serre spectral sequence is multiplicative, so $d_{6}$ satisfies the Leibniz rule
$d_{6}(x\cdot u_{5})=d_{6}(x)\cdot u_{5}+x\cdot d_{6}(u_{5}).$
But $x$ comes from the base and lies in filtration degree $(s,0)$, so $d_{6}(x)=0$ for degree reasons (its target would have negative fibre degree). Hence
$d_{6}(x\cdot u_{5})=x\cdot d_{6}(u_{5}),$
as claimed.
In the present range $s+5\leq 10$ the only nontrivial possibility is $s=4$, yielding a class in bidegree $(4,5)$ and a target in bidegree $(10,0)$. This shows that, within the submodule generated by $u_{5}$, all $d_{r}$ with $2\leq r\leq 5$ vanish and every $d_{6}$ is obtained from $d_{6}(u_{5})$ by multiplication with elements of $H^{*}(BG_{2})$.
Combining this with part (a) completes the proof of the lemma. ∎
Remark 4.3.
  • (i)
    By Lemma 4.1, the class $u=1\otimes u_{5}\in E_{2}^{0,5}(k)$ survives at least to $E_{6}^{0,5}(k)$, and the only potentially nonzero differential with source $u$ in total degree $j\leq 10$ is
    $d_{6}(u)\in E_{6}^{6,0}(k)\cong H^{6}(BG_{2})\otimes M^{0}\cong H^{6}(BG_{2}).$
    Since $H^{6}(BG_{2})\cong\mathbb{F}_{2}\{x_{6}\}$, there exists a uniquely determined scalar $\varepsilon(k)\in\mathbb{F}_{2}$ such that
    $d_{6}(u)=\varepsilon(k)\,x_{6}.$
    This is the scalar appearing in Theorem 1.1(iii).
  • (ii)
    Lemma 4.2 only analyzes the differentials with source in the $H^{*}(BG_{2})$–submodule
    $U\;=\;H^{*}(BG_{2})\cdot u_{5}\;\subset\;E_{2}^{*,*}(k).$
    We do not claim that $M^{j}=H^{j}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$ vanishes for $j\geq 6$, and additional fibre generators in these degrees could in principle support nontrivial differentials in the Serre spectral sequence. Such differentials, however, do not change the description of the possible $d_{6}$–differentials on $U$ in total degree $\leq 10$, which is all that is needed for Theorem 1.1. A determination of $M$ in higher degrees would require an explicit computation via Eilenberg–Moore spectral sequences, and lies beyond the scope of the present paper.
Note 4.4.
The case $k=0$ admits a particularly simple description. The class $0\in\pi_{4}(BG_{2})$ corresponds to the trivial principal $G_{2}$–bundle $S^{4}\times G_{2}\to S^{4}$, and by Gottlieb’s theorem [5], we have
$B\mathcal{G}_{0}\;\simeq\;\mathrm{Map}_{0}(S^{4},BG_{2}),$
where $\mathrm{Map}_{0}(S^{4},BG_{2})$ denotes the component containing the constant map. The evaluation map $\mathrm{ev}\colon\mathrm{Map}_{0}(S^{4},BG_{2})\to BG_{2}$ admits a section $s\colon BG_{2}\to\mathrm{Map}_{0}(S^{4},BG_{2})$ defined by sending $y\in BG_{2}$ to the constant map at $y$. The existence of this section implies that the homomorphism
$\mathrm{ev}^{*}\colon H^{*}(BG_{2})\longrightarrow H^{*}(B\mathcal{G}_{0})$
is injective. In the Serre spectral sequence for the evaluation fibration with $k=0$, the class $x_{6}\in H^{6}(BG_{2})\cong E_{2}^{6,0}$ must therefore survive to $E_{\infty}$ and cannot lie in the image of any differential. In particular, $x_{6}$ cannot be hit by $d_{6}(u_{5})$, forcing $d_{6}(u_{5})=0$. Thus, in our notation, $\varepsilon(0)=0$.
Lemma 4.5 (A 22–local periodicity of $\varepsilon(k)$).
Work 22–locally. If $k\equiv k^{\prime}\pmod{4}$, then
$\varepsilon(k)=\varepsilon(k^{\prime})\in\mathbb{F}_{2}.$
In particular, $\varepsilon(k)$ depends only on the residue class of $k$ modulo $4$.
Proof.
By [8, Lemma 2.1], the connecting map $\partial_{k}\colon G_{2}\to\Omega^{3}_{0}G_{2}$ is the triple adjoint of the Samelson product $\langle k\cdot i_{3},1\rangle$, and in particular $\partial_{k}\simeq k\cdot\partial_{1}$. Kameko proved that the 22–primary component of $\langle i_{3},1\rangle$ has order $4$ [7, Theorem 1.1]. Equivalently, after localization at $2$ the map $\partial_{1}$ has order $4$, so $\partial_{k}\simeq\partial_{k^{\prime}}$ 22–locally whenever $k\equiv k^{\prime}\pmod{4}$.
The Serre spectral sequence (1.1) is functorial with respect to maps of fibrations. Since the kk–dependence of the evaluation fibration is controlled by $\partial_{k}$ (see Section 3), the homotopy $\partial_{k}\simeq\partial_{k^{\prime}}$ induces an identification of the Serre spectral sequences for $k$ and $k^{\prime}$ which is compatible with the standard identifications $E_{2}^{s,t}(k)\cong H^{s}(BG_{2})\otimes H^{t}(\Omega^{3}_{0}G_{2})$. Under this identification, the classes $u_{5}\in E_{2}^{0,5}$ and $x_{6}\in E_{2}^{6,0}\cong H^{6}(BG_{2})$ correspond, and hence the $d_{6}$–values correspond. Since $H^{6}(BG_{2})\cong\mathbb{F}_{2}\{x_{6}\}$ is one-dimensional, the scalar in
$d_{6}(u_{5})=\varepsilon(k)\,x_{6}$
must coincide for $k$ and $k^{\prime}$, proving $\varepsilon(k)=\varepsilon(k^{\prime})$. ∎
Corollary 4.6.
Work 22–locally. If $4\mid k$, then $\varepsilon(k)=0$.
Proof.
If $4\mid k$ then $k\equiv 0\pmod{4}$, so Lemma 4.5 gives $\varepsilon(k)=\varepsilon(0)$. By Note 4.4 we have $\varepsilon(0)=0$, hence $\varepsilon(k)=0$. ∎
Remark 4.7.
Lemma 4.5 and Corollary 4.6 show that, in the low-degree range considered here, the potential kk–dependence through the first fibre class $u_{5}$ reduces to at most three cases:
$k\equiv 0,\ 2,\ 1\text{ or }3\pmod{4}$.
In particular, every class $k\equiv 0\pmod{4}$ falls into the same low-degree behaviour as the trivial bundle, namely $\varepsilon(k)=0$. Thus, to detect genuinely new behaviour via $d_{6}(u_{5})$, it suffices to determine $\varepsilon(1)$ and $\varepsilon(2)$. This reduction is consistent with the 22–local classification of $G_{2}$–gauge groups governed by the 22–primary order of $\langle i_{3},1\rangle$ [8, 7].
4.3. The $E_{\infty}$–page and completion of the proof
We now describe the $E_{\infty}$–page in total degree $\leq 10$. By Lemma 4.2, the only potentially nonzero differential with source or target in this range on the submodule generated by $u_{5}$ is $d_{6}$ on $E_{6}^{0,5}(k)$. All other classes in total degree $\leq 10$ are permanent cycles.
Let us first consider the case $\varepsilon(k)=0$. Then $d_{6}(u)=0$, and by multiplicativity $d_{6}=0$ on all of $E_{6}^{*,*}(k)$ in total degree $\leq 10$. It follows that $E_{2}^{*,*}(k)\cong E_{\infty}^{*,*}(k)$ in this range, and the Serre spectral sequence collapses at $E_{2}$ in total degree $\leq 10$.
If $\varepsilon(k)=1$, then $d_{6}(u)=x_{6}$, and the class $x_{6}$ does not survive to $E_{\infty}^{6,0}(k)$. More generally, the image of $d_{6}$ is the $H^{*}(BG_{2})$–submodule generated by $x_{6}$, and $E_{\infty}^{*,*}(k)^{\leq 10}$ is obtained from $H^{*}(BG_{2})\otimes M^{\leq 10}$ by imposing the single relation $x_{6}=d_{6}(u_{5})$ in total degree $\leq 10$.
In either case, for each total degree $j\leq 10$ there is a finite filtration of $H^{j}(B\mathcal{G}_{k})$ whose associated graded pieces are the groups $E_{\infty}^{s,t}(k)$ with $s+t=j$. This gives the short exact sequence of Theorem 1.1(v), with $Q_{j}(k)$ recording possible extension phenomena between different bidegrees. Since the only dependence on $k$ in the spectral sequence occurs via the scalar $\varepsilon(k)$, it follows that all potential kk–dependence in $H^{j}(B\mathcal{G}_{k})$ for $j\leq 10$ is encoded in this single differential.
Proof of Theorem 1.1.
Part (i) is the identification of the $E_{2}$–term with trivial local coefficients. Part (ii) is Proposition 3.2, and parts (iii) and (iv) follow from Lemmas 4.1 and 4.2, together with the discussion above of the differential $d_{6}(u)=\varepsilon(k)x_{6}$. Part (v) is the standard description of $H^{*}(B\mathcal{G}_{k})$ in terms of the $E_{\infty}$–term and the resulting finite filtration. This completes the proof. ∎
Remark 4.8.
The restriction to total degree $j\leq 10$ in Theorem 1.1 is motivated by two factors. First, this range is sufficient to capture the first nontrivial cohomological dependence on the bundle class $k$, which manifests via the differential $d_{6}(u_{5})$. Second, extending the analysis beyond degree $10$ requires a substantially more detailed understanding of the fibre cohomology $M=H^{*}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$ as an unstable $\mathcal{A}$–module, which involves Eilenberg–Moore spectral sequence computations beyond the scope of this paper. Once $M$ is explicitly determined in a larger range, the computation of $H^{*}(B\mathcal{G}_{k})$ reduces to a finite problem in the “hit” theory over the polynomial algebra $H^{*}(BG_{2})$, well suited to computer implementation in systems such as SageMath. We hope to return to these global computations in future work.