### v4

# Low-degree mod 2 cohomology of classifying spaces of $G_{2}$-gauge groups

Abstract.
Let $G$ be a simply connected compact simple Lie group and let $\mathcal{G}_{k}$ denote the gauge group of a principal GG–bundle over $S^{4}$ with second Chern class $k\in\pi_{4}(BG)\cong\mathbb{Z}$. For $G=G_{2}$, the pp–local homotopy types of the gauge groups have been completely classified through the work of Kishimoto–Theriault–Tsutaya and Kameko, in terms of the order of the fundamental Samelson product $\langle i_{3},1\rangle\in[\Sigma^{3}G_{2},G_{2}]$.
In this paper, we begin a complementary study of the mod $2$ cohomology of the classifying spaces $B\mathcal{G}_{k}(G_{2})$. Our goal is to understand the structure of $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ as an unstable module over the mod $2$ Steenrod algebra in a low range of degrees. Using the evaluation fibration
$\Omega_{0}^{3}G_{2}\longrightarrow B\mathcal{G}_{k}\xrightarrow{\;\mathrm{ev}\;}BG_{2}$
and the associated Serre spectral sequence, together with low-degree information on $\Omega^{3}_{0}G_{2}$, we study
$H^{s}(BG_{2};H^{t}(\Omega^{3}_{0}G_{2}))\Longrightarrow H^{s+t}(B\mathcal{G}_{k})$
in total degree $\leq 10$. A careful analysis of the homotopy groups of $G_{2}$ shows that
$H^{j}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})=0\quad\text{for }1\leq j\leq 4,\qquad H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})\cong\mathbb{F}_{2},$
so the least positive degree in which the fibre cohomology is nonzero is $5$. As a consequence, there is a distinguished generator
$u_{5}\in H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$
such that, in total degree $\leq 10$, the only possible Serre differential with source $u_{5}$ is a $d_{6}$–differential
$d_{6}(u_{5})=\varepsilon(k)\,x_{6},$
where $x_{6}\in H^{6}(BG_{2};\mathbb{F}_{2})$ is the degree-$6$ generator and $\varepsilon(k)\in\mathbb{F}_{2}$ records the contribution of the bundle class to this first nontrivial fibre differential. In addition, 22–locally we prove that $\varepsilon(k)$ is 88–periodic in $k$ (i.e. it depends only on $k\bmod 8$) and that $\varepsilon(k)=0$ for all $k\equiv 0\pmod{8}$.
Key words and phrases:
Gauge groups, Classifying spaces, Steenrod algebra, Serre spectral sequence
2020 Mathematics Subject Classification:
Primary 55R35; Secondary 55S10, 55P35, 55T10, 55T20
Acknowledgment
The author would like to express his sincere gratitude to the anonymous referee for the careful reading of the manuscript and for the valuable suggestions and corrections, which have greatly improved the clarity and overall quality of the paper.
1. Introduction
1.1. Gauge groups and their homotopy types
Let $G$ be a compact, connected Lie group, let $X$ be a pointed CW complex, and let $P\to X$ be a principal GG–bundle. The gauge group $\mathcal{G}(P)$ is the topological group of GG–equivariant bundle automorphisms of $P$ that cover the identity on $X$. When $X$ is finite, Crabb–Sutherland showed that there are only finitely many homotopy types among the gauge groups $\mathcal{G}(P)$ as $P$ ranges over all principal GG–bundles over $X$ [6]. In the fundamental case $X=S^{4}$ and $G$ simply connected and simple, isomorphism classes of GG–bundles are classified by $\pi_{4}(BG)\cong\mathbb{Z}$, and we write $\mathcal{G}_{k}(G)$ for the gauge group of the bundle classified by $k\in\mathbb{Z}$.
The problem of determining the homotopy types of these gauge groups for specific Lie groups $G$ has received sustained attention; see for example [11, 12, 16]. In many cases the classification is governed by the order of a fundamental Samelson product in $G$.
The case $G=G_{2}$ is particularly interesting because of the exceptional nature of $G_{2}$. Kishimoto, Theriault and Tsutaya  [11] proved that the classification for $G_{2}$–gauge groups reduces to determining the order of the Samelson product
$\langle i_{3},1\rangle\in[\Sigma^{3}G_{2},G_{2}],$
where $i_{3}\colon S^{3}\to G_{2}$ is the inclusion of the bottom cell and $1$ denotes the identity map. Moreover, their work determines the number of homotopy types up to one factor of $2$ [11]. Kameko subsequently resolved the remaining 22–primary ambiguity by proving that the order of $\langle i_{3},1\rangle$ is exactly $168$ [10, Theorem 1.4]. Consequently, the pp–local homotopy classification of $\mathcal{G}_{k}(G_{2})$ is determined by the greatest common divisor $(k,168)$; at the prime $2$, it is determined by $(k,8)$ [10, Theorem 1.1 and Proposition 1.2]. Thus the homotopy types of the gauge groups are completely understood up to pp–localization. It is natural to ask how far this information is reflected in the (co)homology of their classifying spaces.
1.2. Cohomology of $B\mathcal{G}_{k}$ and the Steenrod algebra
Gottlieb identified the classifying space of $\mathcal{G}_{k}$ with a component of a mapping space [7]: specifically, $B\mathcal{G}_{k}$ is homotopy equivalent to $\mathrm{Map}_{k}(S^{4},BG)$, the path component of $\mathrm{Map}(S^{4},BG)$ containing a classifying map representing $k\in\pi_{4}(BG)\cong\mathbb{Z}$. This identification yields a fundamental evaluation fibration
$\Omega^{3}_{0}G\;\longrightarrow\;B\mathcal{G}_{k}\xrightarrow{\;\mathrm{ev}\;}BG.$
For $G=G_{2}$, this expresses $B\mathcal{G}_{k}$ as a twisted mapping space over the base $BG_{2}$.
The mod $2$ cohomology of $BG_{2}$ is classically known: Borel showed that
$H^{*}(BG_{2};\mathbb{F}_{2})\cong\mathbb{F}_{2}[x_{4},x_{6},x_{7}]$
is a polynomial algebra on generators in degrees $4,6,7$ [2]. The action of the Steenrod algebra $\mathcal{A}$ on $H^{*}(BG_{2})$ can be described explicitly via Wu formulas, and from the viewpoint of the hit problem one is interested in minimal generating sets for the unstable $\mathcal{A}$–module $\mathbb{F}_{2}[x_{4},x_{6},x_{7}]$.
By contrast, there seems to be no published description of the full mod $2$ cohomology ring $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ for nontrivial $G_{2}$–bundles over $S^{4}$, nor of its structure as an unstable $\mathcal{A}$–module. There are, however, important partial precedents. Choi studied the mod $p$ homology of classifying spaces of $\mathrm{Sp}(n)$–gauge groups over $S^{4}$ [4], and he also computed the homology of the $G_{2}$–gauge groups themselves using Eilenberg–Moore and Serre spectral sequences [5]. Our aim here is to give a precise structural description of $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ in low degrees, in terms of the evaluation fibration and the first nontrivial cohomology of the triple loop space $\Omega^{3}_{0}G_{2}$.
1.3. Main result: a low-degree description
We work 22–locally throughout and write $H^{*}(-)$ for $H^{*}(-;\mathbb{F}_{2})$. Consider the Serre spectral sequence of the evaluation fibration
(1.1) $E_{2}^{s,t}(k)\;\cong\;H^{s}(BG_{2};H^{t}(\Omega^{3}_{0}G_{2}))\;\Longrightarrow\;H^{s+t}(B\mathcal{G}_{k}).$
The $E_{2}$–term is a module over $H^{*}(BG_{2})$ and an unstable module over the Steenrod algebra. The twisting of this spectral sequence depends on $k$ through the connecting map
$\partial_{k}\colon G_{2}\longrightarrow\Omega^{3}_{0}G_{2},$
which, by Lang’s description of the boundary map of the evaluation fibration and its formulation in the gauge-group setting [11, 13], is the triple adjoint of the Samelson product $\langle k\cdot i_{3},1\rangle$. Kameko’s calculation that the 22–primary order of $\langle i_{3},1\rangle$ is $8$ implies that, after localization at $2$, the map $\partial_{k}$ depends only on $k\bmod 8$ [10, Proposition 1.3].
The cohomology of the fibre $\Omega^{3}_{0}G$ can in principle be computed by iterated Eilenberg–Moore spectral sequences applied to the path-loop fibrations of $BG$ and $G$. Choi and Yoon conjectured that, for any simply connected finite HH–space $X$, the Eilenberg–Moore spectral sequences converging to the mod $p$ homology of $\Omega^{2}X$ and $\Omega^{3}X$ collapse at $E_{2}$ [3]. This conjecture was subsequently proved by Lin [14]; in particular, for any compact simple Lie group $G$ and any prime $p$, the Eilenberg–Moore spectral sequences computing $H_{*}(\Omega^{2}G;\mathbb{F}_{p})$ and $H_{*}(\Omega^{3}G;\mathbb{F}_{p})$ collapse at $E_{2}$. Combined with the classical description of $H^{*}(BG_{2})$, this implies that $H^{*}(\Omega^{3}_{0}G_{2})$ is of finite type in each degree and that its generators occur in bounded degrees. For our purposes, however, we only need to know the very first nontrivial cohomology group of $\Omega^{3}_{0}G_{2}$.
Let
$M\;=\;H^{*}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$
and write $M^{\leq N}=\bigoplus_{j\leq N}M^{j}$ for its truncation in degrees $\leq N$. Our main theorem describes the Serre spectral sequence (1.1) in total degree $\leq 10$ and singles out a distinguished scalar $\varepsilon(k)\in\mathbb{F}_{2}$ which already captures the effect of $k$ on the first nontrivial fibre class.
Theorem 1.1 (Low-degree structure of $H^{*}(B\mathcal{G}_{k})$).
Work 22–locally and fix an integer $k$. Consider the Serre spectral sequence (1.1) for the fibration $\Omega^{3}_{0}G_{2}\to B\mathcal{G}_{k}\to BG_{2}$, and let $M=H^{*}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$.
  1. (i)
    In total degree $j=s+t\leq 10$ the $E_{2}$–term is naturally isomorphic, as a bigraded $H^{*}(BG_{2})$–module, to
    $E_{2}^{*,*}(k)^{\leq 10}\;\cong\;\bigl(H^{*}(BG_{2})\otimes M\bigr)^{\leq 10},$
    where the superscript ≤10 denotes truncation in total degree. Equivalently, $E_{2}^{s,t}(k)\cong H^{s}(BG_{2})\otimes M^{t}$ for $s+t\leq 10$.
  2. (ii)
    Using the low-degree homotopy groups of $G_{2}$ together with integral and 22–local Hurewicz arguments, one has
    $M^{j}=H^{j}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})=0\quad\text{for }1\leq j\leq 4,\qquad M^{5}\cong\mathbb{F}_{2}.$
    In particular, there exists a generator
    $u_{5}\in M^{5}\cong E_{2}^{0,5}(k)$
    and $u_{5}$ is a generator of least positive degree in $H^{*}(\Omega^{3}_{0}G_{2})$.
  3. (iii)
    In total degree $j\leq 10$, all differentials
    $d_{r}(u_{5})\colon E_{r}^{0,5}(k)\longrightarrow E_{r}^{r,\,5-r+1}(k)$
    vanish for $2\leq r\leq 5$ and for $r\geq 7$, and the only page on which $u_{5}$ can support a nontrivial differential is $r=6$. After choosing a generator $x_{6}\in H^{6}(BG_{2})\cong E_{2}^{6,0}(k)$ and the generator $u_{5}\in M^{5}\cong E_{2}^{0,5}(k)$, there is a uniquely determined scalar $\varepsilon(k)\in\mathbb{F}_{2}$ such that
    $d_{6}(u_{5})=\varepsilon(k)\,x_{6}$
    in $E_{6}^{6,0}(k)$.
  4. (iv)
    Let $U\subset E_{2}^{*,*}(k)$ denote the $H^{*}(BG_{2})$–submodule generated by $u_{5}$. In total degree $j\leq 10$ all differentials $d_{r}$ with $2\leq r\leq 5$ vanish on $U$, and every differential
    $d_{6}\colon E_{6}^{s,5}(k)\longrightarrow E_{6}^{s+6,0}(k)$
    with $s+5\leq 10$ satisfies
    $d_{6}(x\cdot u_{5})=x\cdot d_{6}(u_{5})\qquad\text{for }x\in H^{s}(BG_{2}).$
    In particular, within $U$ the entire effect of the Serre differentials in total degree $\leq 10$ is determined by the single scalar $\varepsilon(k)\in\mathbb{F}_{2}$.
  5. (v)
    For each $j\leq 10$, the spectral sequence induces a finite filtration
    $0=F^{j+1}H^{j}(B\mathcal{G}_{k})\subseteq F^{j}H^{j}(B\mathcal{G}_{k})\subseteq\cdots\subseteq F^{0}H^{j}(B\mathcal{G}_{k})=H^{j}(B\mathcal{G}_{k})$
    whose associated graded object is
    $\mathrm{gr}\,H^{j}(B\mathcal{G}_{k})\cong\bigoplus_{s+t=j}E_{\infty}^{s,t}(k).$
    In total degree $j\leq 10$, the contribution to $E_{\infty}^{*,*}(k)$ coming from the $H^{*}(BG_{2})$–submodule generated by $u_{5}$ is explicitly determined by the single scalar $\varepsilon(k)\in\mathbb{F}_{2}$.
In particular, in total degree $j\leq 10$ the dependence of $H^{j}(B\mathcal{G}_{k};\mathbb{F}_{2})$ on the bundle class $k$ that is visible through the first nontrivial fibre class $u_{5}$ is completely encoded in the single Serre differential $d_{6}(u_{5})=\varepsilon(k)x_{6}$.
Remark 1.2.
Theorem 1.1 should be viewed as a structural reduction. It does not claim a closed formula for $\varepsilon(k)$, nor does it assert that $\varepsilon(k)$ is nontrivial. Kameko’s calculation shows that the 22–primary order of $\langle i_{3},1\rangle$ is $8$, while the 22–local homotopy classification of the gauge groups is indexed by $(k,8)$ [10, Proposition 1.2 and Proposition 1.3]. It is therefore natural to expect that $\varepsilon(k)$ is governed by the same 22–primary data.
In fact, we prove (Lemma 4.4) that, 22–locally, the scalar $\varepsilon(k)$ depends only on the residue class of $k$ modulo $8$, and moreover $\varepsilon(k)=0$ whenever $8\mid k$ (Corollary 4.5). Thus the argument given here reduces the remaining low-degree problem to determining the seven values $\varepsilon(1),\ldots,\varepsilon(7)$. Although Kameko’s classification identifies the 22–local homotopy types of the gauge groups by the invariant $(k,8)$, the proof below uses the stronger condition that the connecting maps agree, namely congruence modulo $8$. Any further reduction among these residue classes would require an additional fibrewise comparison argument, which we do not establish here.
Determining these values explicitly would require a more detailed understanding of $H^{*}(\Omega^{3}_{0}G_{2})$ and of the induced map of the connecting map $\partial_{k}$ on mod $2$ cohomology, and we leave this as a natural problem for future work.
Remark 1.3 (Applications and significance).
Although Theorem 1.1 is a low-degree result, it has several concrete applications.
First, it provides a first cohomological manifestation of the dependence of the classifying spaces $B\mathcal{G}_{k}$ on the bundle class $k$, through the first nontrivial fibre differential. In this sense, the scalar $\varepsilon(k)\in\mathbb{F}_{2}$ may be viewed as a low-degree cohomological shadow of the 22–local homotopy classification of the gauge groups $\mathcal{G}_{k}(G_{2})$.
Second, the theorem reduces the first kk–dependent part of the low-degree analysis of $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ to two more concrete problems: determining the next stages of the fibre cohomology $H^{*}(\Omega_{0}^{3}G_{2};\mathbb{F}_{2})$ beyond degree $5$, and determining the values of the scalar $\varepsilon(k)$. Thus the remaining difficulty becomes much more explicit and the theorem provides a practical framework for further calculations.
Third, since our goal is to study $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ as an unstable module over the Steenrod algebra, the present low-degree description gives the first structural input for any future analysis of generators, Steenrod operations, and hit-theoretic questions for these classifying spaces.
Finally, because $\mathcal{G}_{k}\simeq\Omega B\mathcal{G}_{k}$, the low-degree information obtained here is also relevant to comparisons with the mod $2$ homology of the gauge groups themselves, and may serve as a starting point for further work relating the classifying-space approach of the present paper to the Eilenberg–Moore and Serre spectral sequence calculations of Choi [5].
In particular, the theorem isolates the first possible obstruction to the survival of the degree-$6$ base class in the evaluation Serre spectral sequence, and thus identifies the first place where the bundle class can affect the mod $2$ cohomology of $B\mathcal{G}_{k}$ through the first nontrivial fibre class.
The paper is organised as follows. In Section 2, we review the necessary background on gauge groups, Samelson products, the mod $2$ cohomology of $BG_{2}$ and the Steenrod algebra. In Section 3, we describe the evaluation fibration and its Serre spectral sequence, and we recall the relationship between the connecting map and the Samelson product. Finally, Section 4 contains the proof of Theorem 1.1, including a detailed analysis of the bidegrees for $j\leq 10$ and the vanishing of $H^{j}(\Omega^{3}_{0}G_{2})$ for $1\leq j\leq 4$.
2. Preliminaries
2.1. Gauge groups and mapping spaces
Let $G$ be a compact, connected Lie group and $P\to S^{4}$ a principal GG–bundle. The associated gauge group $\mathcal{G}(P)$ is the group of GG–equivariant automorphisms of $P$ covering the identity on $S^{4}$, with the compact-open topology. When $G$ is simply connected and simple, its classifying space $BG$ is 33–connected and $\pi_{4}(BG)\cong\mathbb{Z}$; the integer $k\in\mathbb{Z}$ corresponding to $[P]$ plays the role of the second Chern class.
Following the notation of [10, 11], we write $\mathcal{G}_{k}=\mathcal{G}_{k}(G_{2})$ for the gauge group of the principal $G_{2}$–bundle classified by $k\in\pi_{4}(BG_{2})\cong\mathbb{Z}$. By a classical result of Gottlieb [7], the classifying space $B\mathcal{G}_{k}$ is homotopy equivalent to $\mathrm{Map}_{k}(S^{4},BG_{2})$, the connected component of the mapping space $\mathrm{Map}(S^{4},BG_{2})$ corresponding to the class $k\in\pi_{4}(BG_{2})\cong\mathbb{Z}$.
Under this identification, the evaluation map $\mathrm{ev}\colon\mathrm{Map}(S^{4},BG_{2})\to BG_{2}$ restricts to a fibration
$\mathrm{Map}_{*}(S^{4},BG_{2})_{k}\longrightarrow B\mathcal{G}_{k}\xrightarrow{\;\mathrm{ev}\;}BG_{2},$
where $\mathrm{Map}_{*}(S^{4},BG_{2})_{k}$ denotes the path component of the based mapping space $\mathrm{Map}_{*}(S^{4},BG_{2})$ corresponding to $k\in\pi_{4}(BG_{2})\cong\mathbb{Z}$. For simply connected $G$, this fibre is homotopy equivalent to $\Omega_{0}^{3}G$.
When $G$ is simply connected, the based mapping space $\mathrm{Map}_{*}(S^{4},BG)_{k}$ is homotopy equivalent to the triple loop space $\Omega^{3}_{0}G$ for every $k$, and via Gottlieb’s equivalence we obtain the fibration
$\Omega^{3}_{0}G\;\longrightarrow\;B\mathcal{G}_{k}(G)\xrightarrow{\;\mathrm{ev}\;}BG.$
We henceforth specialise to $G=G_{2}$.
2.2. Samelson products and the connecting map
Let $G$ be an HH–space. The Samelson product
$\langle-,-\rangle\colon\pi_{m}(G)\otimes\pi_{n}(G)\longrightarrow\pi_{m+n}(G)$
is defined by the composite
$S^{m}\wedge S^{n}\xrightarrow{f\wedge g}G\wedge G\xrightarrow{[-,-]}G,$
where $f,g$ represent the chosen homotopy classes and [,][-,-] is the commutator in $G$. For a compact simple Lie group $G$ the order of certain Samelson products controls the classification of gauge groups; see for example [11, 12, 16] for concrete instances of this phenomenon.
In the present context the relevant Samelson product is
$\langle i_{3},1\rangle\in[\Sigma^{3}G,G],$
where $i_{3}\colon S^{3}\to G$ is the inclusion of the bottom cell and $1$ is the identity. Lang [13] showed that the boundary map of the evaluation fibration is identified with the adjoint of a Whitehead product. In the present gauge-group setting, this implies that the connecting map
$\partial_{k}\colon G\longrightarrow\Omega^{3}_{0}G$
in the homotopy fibration
$\mathcal{G}_{k}(G)\;\longrightarrow\;G\xrightarrow{\;\partial_{k}\;}\Omega^{3}_{0}G\longrightarrow B\mathcal{G}_{k}(G)\xrightarrow{\;\mathrm{ev}\;}BG$
is the triple adjoint of the Samelson product $\langle k\cdot i_{3},1\rangle$; see [11, Lemma 2.1] for the formulation used here. In particular $\partial_{k}\simeq k\cdot\partial_{1}$, and the order of $\partial_{1}$ is equal to the order of $\langle i_{3},1\rangle$.
For $G=G_{2}$, Kameko’s corrected calculation [10, Proposition 1.3] implies that the 22–primary component of $\langle i_{3},1\rangle$ has order $8$. Thus, after localizing at $2$, the homotopy class of $\partial_{k}$ depends only on $k\bmod 8$. This 22–primary information is the only input from the classification of $G_{2}$–gauge groups that will enter our cohomological analysis.
3. The evaluation fibration and its spectral sequence
3.1. The Serre spectral sequence
Specialising the evaluation fibration to $G=G_{2}$, we obtain a fibration of simply connected spaces
$\Omega^{3}_{0}G_{2}\;\longrightarrow\;B\mathcal{G}_{k}\xrightarrow{\;\mathrm{ev}\;}BG_{2}$
and hence a mod $2$ cohomology Serre spectral sequence
$E_{2}^{s,t}(k)=H^{s}(BG_{2};H^{t}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2}))\;\Longrightarrow\;H^{s+t}(B\mathcal{G}_{k};\mathbb{F}_{2}).$
We abbreviate $M=H^{*}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$; it is a graded commutative algebra and an unstable $\mathcal{A}$–module.
Since $BG_{2}$ is simply connected and the action of $\pi_{1}(BG_{2})$ on $H^{*}(\Omega^{3}_{0}G_{2})$ is trivial, the local coefficient system on the fibre is trivial, and we have
$E_{2}^{s,t}(k)\cong H^{s}(BG_{2})\otimes M^{t}$
as bigraded vector spaces, with multiplicative structure induced from the tensor product.
The spectral sequence is natural with respect to maps of fibrations. In particular, the connecting map $\partial_{k}\colon G_{2}\to\Omega^{3}_{0}G_{2}$ in the associated homotopy fibration controls the kk–dependence of the differentials.
3.2. The triple loop space $\Omega^{3}_{0}G_{2}$
The homotopy groups of $G_{2}$ in low degrees were computed by Mimura and Toda [15]. For our purposes we only need the following facts:
$\pi_{3}(G_{2})\cong\mathbb{Z},\qquad\pi_{4}(G_{2})=0,\qquad\pi_{5}(G_{2})=0,\qquad\pi_{6}(G_{2})\text{ is $3$--torsion},\qquad\pi_{7}(G_{2})=0,$
and $\pi_{8}(G_{2})$ contains 22–torsion. The precise structure of $\pi_{6}(G_{2})$ and $\pi_{8}(G_{2})$ will not be needed.
We focus on the component $\Omega^{3}_{0}G_{2}$ containing the constant map.
Lemma 3.1.
The space $\Omega^{3}_{0}G_{2}$ is 22–connected. In particular
$\pi_{1}(\Omega^{3}_{0}G_{2})=\pi_{2}(\Omega^{3}_{0}G_{2})=0.$
Proof.
By definition, $\Omega^{3}_{0}G_{2}$ denotes the path component of the constant map in the triple loop space $\Omega^{3}G_{2}$, so $\pi_{i}(\Omega^{3}_{0}G_{2})\cong\pi_{i}(\Omega^{3}G_{2})$ for all $i\geq 1$ and $\pi_{0}(\Omega^{3}_{0}G_{2})=0$.
For any based space $X$ and integer $n\geq 0$, there is a natural isomorphism
$\pi_{n}(\Omega X)\;\cong\;\pi_{n+1}(X),$
obtained by identifying both sides with homotopy classes of based maps via the suspension loop adjunction
[…]
Iterating this identification three times gives, for every $i\geq 0$,
$\pi_{i}(\Omega^{3}G_{2})\;\cong\;\pi_{i+1}(\Omega^{2}G_{2})\;\cong\;\pi_{i+2}(\Omega G_{2})\;\cong\;\pi_{i+3}(G_{2}).$
Since $\Omega^{3}_{0}G_{2}$ is the identity component of $\Omega^{3}G_{2}$, we in particular obtain
$\pi_{i}(\Omega^{3}_{0}G_{2})\;\cong\;\pi_{i+3}(G_{2})\qquad\text{for }i\geq 1.$
The low-degree homotopy groups of $G_{2}$ were computed by Mimura–Toda [15]; in particular
$\pi_{4}(G_{2})=0,\qquad\pi_{5}(G_{2})=0.$
Substituting $i=1,2$ in the isomorphism above yields
$\pi_{1}(\Omega^{3}_{0}G_{2})\cong\pi_{4}(G_{2})=0,\qquad\pi_{2}(\Omega^{3}_{0}G_{2})\cong\pi_{5}(G_{2})=0.$
Thus $\Omega^{3}_{0}G_{2}$ is 22–connected, as claimed. ∎
We can now determine the first few homology and cohomology groups of $\Omega^{3}_{0}G_{2}$ using the Hurewicz theorem and universal coefficients.
Proposition 3.2.
Let $M=H^{*}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$. Then:
  1. (a)
    $M^{0}\cong\mathbb{F}_{2}$ and
    $M^{1}=M^{2}=M^{3}=M^{4}=0.$
  2. (b)
    One has
    $M^{5}=H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})\cong\mathbb{F}_{2}.$
    In particular, if $0\neq u_{5}\in M^{5}$, then $u_{5}$ is a generator of $M^{5}$ and is of least positive degree in $H^{*}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$.
Proof.
Let $X=\Omega^{3}_{0}G_{2}$. By Lemma 3.1, $X$ is 22–connected.
Step 1: integral information in degrees $\leq 4$. Since $X$ is 22–connected, the Hurewicz theorem gives
$H_{1}(X;\mathbb{Z})=H_{2}(X;\mathbb{Z})=0,$
an isomorphism
$h\colon\pi_{3}(X)\xrightarrow{\cong}H_{3}(X;\mathbb{Z}),$
and a surjection
[…]
Using $\pi_{i}(X)\cong\pi_{i+3}(G_{2})$ for $i\geq 1$ and the low-degree homotopy groups of $G_{2}$, we obtain
$\pi_{3}(X)\cong\pi_{6}(G_{2})\quad\text{(a finite $3$--group)},\qquad\pi_{4}(X)\cong\pi_{7}(G_{2})=0.$
Hence
$H_{3}(X;\mathbb{Z})\cong\pi_{3}(X)\cong\pi_{6}(G_{2})$
is a finite 33–group, while
$H_{4}(X;\mathbb{Z})=0.$
Now apply the universal coefficient short exact sequence for cohomology. Since $H_{1}(X;\mathbb{Z})=H_{2}(X;\mathbb{Z})=0$, we have
$H^{1}(X;\mathbb{F}_{2})=H^{2}(X;\mathbb{F}_{2})=0.$
Since $H_{3}(X;\mathbb{Z})$ is a finite 33–group, both
$\operatorname{Hom}_{\mathbb{Z}}(H_{3}(X;\mathbb{Z}),\mathbb{F}_{2})=0\qquad\text{and}\qquad\operatorname{Ext}^{1}_{\mathbb{Z}}(H_{3}(X;\mathbb{Z}),\mathbb{F}_{2})=0.$
Using also $H_{4}(X;\mathbb{Z})=0$, it follows that
$H^{3}(X;\mathbb{F}_{2})=H^{4}(X;\mathbb{F}_{2})=0.$
This proves part (a).
Step 2: the first nonzero class in degree $5$. We now work 22–locally. Since $X$ is nilpotent, localization at $2$ induces an isomorphism on mod $2$ homology
[…]
Moreover, for each $i\geq 1$ one has
$\pi_{i}(X_{(2)})\cong\pi_{i}(X)\otimes\mathbb{Z}_{(2)}.$
Because
$\pi_{1}(X)=\pi_{2}(X)=0,\qquad\pi_{3}(X)\cong\pi_{6}(G_{2})\ \text{is $3$--primary},\qquad\pi_{4}(X)\cong\pi_{7}(G_{2})=0,$
it follows that
$\pi_{i}(X_{(2)})=0\qquad\text{for }1\leq i\leq 4.$
Thus $X_{(2)}$ is 44–connected.
Applying the integral Hurewicz theorem to the 44–connected space $X_{(2)}$, we obtain an isomorphism
[…]
Using again $\pi_{i}(X)\cong\pi_{i+3}(G_{2})$ together with the classical computation
$\pi_{8}(G_{2})\cong\mathbb{Z}/2,$
we get
$\pi_{5}(X_{(2)})\cong\pi_{5}(X)\otimes\mathbb{Z}_{(2)}\cong\pi_{8}(G_{2})\otimes\mathbb{Z}_{(2)}\cong\mathbb{Z}/2.$
Hence
$H_{5}(X_{(2)};\mathbb{Z})\cong\mathbb{Z}/2.$
Reducing coefficients mod $2$ gives
$H_{5}(X_{(2)};\mathbb{F}_{2})\cong\mathbb{F}_{2}.$
Using invariance of mod $2$ homology under 22–localization, we conclude that
$H_{5}(X;\mathbb{F}_{2})\cong\mathbb{F}_{2}.$
Since coefficients are taken in the field $\mathbb{F}_{2}$, the universal coefficient theorem gives
$H^{5}(X;\mathbb{F}_{2})\cong\operatorname{Hom}_{\mathbb{F}_{2}}(H_{5}(X;\mathbb{F}_{2}),\mathbb{F}_{2})\cong\mathbb{F}_{2}.$
This proves part (b). ∎
For the remainder of the paper, we fix a generator
$u_{5}\in M^{5}=H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})\cong\mathbb{F}_{2}.$
4. Proof of the low-degree theorem
In this section we analyse the Serre spectral sequence (1.1) in total degree $j=s+t\leq 10$ and prove Theorem 1.1. We work throughout with mod $2$ coefficients and suppress them from the notation.
4.1. The $E_{2}$–page in low degrees
By the discussion in Section 3, the $E_{2}$–term has the form
$E_{2}^{s,t}(k)\cong H^{s}(BG_{2})\otimes M^{t},$
and $H^{*}(BG_{2})\cong\mathbb{F}_{2}[x_{4},x_{6},x_{7}]$ is generated by $x_{4},x_{6},x_{7}$ in degrees $4,6,7$. For $0<j\leq 10$ the only degrees $s>0$ in which $H^{s}(BG_{2})$ can be nonzero are
$s\in\{4,6,7,8,10\},$
corresponding to monomials $x_{4}$, $x_{6}$, $x_{7}$, $x_{4}^{2}$ and $x_{4}x_{6}$, respectively.
By Proposition 3.2, the fibre cohomology $M^{t}$ vanishes for $1\leq t\leq 4$. Consequently, in total degree $j=s+t\leq 10$ the only nonzero groups $E_{2}^{s,t}$ with $t>0$ occur when $t\geq 5.$ In particular, the class $u_{5}\in M^{5}$ gives rise to a class
$u=1\otimes u_{5}\in E_{2}^{0,5}(k),$
and this is a generator of least positive degree in $M$.
4.2. Bidegree constraints on differentials
The differentials in the Serre spectral sequence have bidegree
$d_{r}\colon E_{r}^{s,t}\longrightarrow E_{r}^{s+r,t-r+1}.$
We now examine which values of $r$ can occur with source or target in total degree $j\leq 10$.
Let $(s,t)$ be a bidegree with $s+t\leq 10$ and $E_{r}^{s,t}\neq 0$. If $d_{r}$ is nonzero on $E_{r}^{s,t}$, then its target lies in bidegree (s+r,tr+1)(s+r,t-r+1) with total degree
$(s+r)+(t-r+1)=s+t+1\leq 11.$
Moreover, for $s>0$ we know that $H^{s}(BG_{2})$ can only be nonzero in degrees
$s\in\{4,6,7,8,10\},$
and by Proposition 3.2 the fibre cohomology satisfies
$M^{0}\cong\mathbb{F}_{2},\qquad M^{j}=0\ \text{for}\ 1\leq j\leq 4,\qquad M^{5}\neq 0.$
In this subsection we restrict attention to differentials whose source is the first nontrivial fibre class
$u_{5}\in M^{5}\cong E_{2}^{0,5}(k),$
and we determine on which page $u_{5}$ can possibly support a nontrivial differential in total degree $j\leq 10$.
Since $u_{5}$ lies in bidegree $(0,5)$, any differential with source $u_{5}$ has the form
$d_{r}(u_{5})\in E_{r}^{r,\,5-r+1}.$
We analyse the possible values of $r$.
  • If $r\geq 7$, then $5-r+1\leq-1$, so the target bidegree has negative fibre degree and $E_{r}^{r,5-r+1}=0$. Thus $d_{r}(u_{5})=0$ for all $r\geq 7$ in total degree $\leq 10$.
  • For $2\leq r\leq 5$ we have the following possibilities:
    • $r=2$: the target is $(2,4)$, but $H^{2}(BG_{2})=0$, so $E_{2}^{2,4}=0$ and hence $d_{2}(u_{5})=0$.
    • $r=3$: the target is $(3,3)$, and again $H^{3}(BG_{2})=0$, so $d_{3}(u_{5})=0$.
    • $r=4$: the target is $(4,2)$; here $H^{4}(BG_{2})\cong\mathbb{F}_{2}\{x_{4}\}$ is nonzero, but $M^{2}=0$ by Proposition 3.2, so $E_{2}^{4,2}\cong H^{4}(BG_{2})\otimes M^{2}=0$, and hence $d_{4}(u_{5})=0$.
    • $r=5$: the target is $(5,1)$, and $H^{5}(BG_{2})=0$, so $E_{2}^{5,1}=0$ and $d_{5}(u_{5})=0$.
    In particular, $d_{r}(u_{5})=0$ for all $2\leq r\leq 5$.
  • For $r=6$ the target is bidegree $(6,0)$, and
    $E_{2}^{6,0}\cong H^{6}(BG_{2})\otimes M^{0}\cong H^{6}(BG_{2})\cong\mathbb{F}_{2}\{x_{6}\}$
    is nonzero. There is therefore no a priori bidegree reason to force $d_{6}(u_{5})$ to vanish; this is the first page on which $u_{5}$ can possibly support a nontrivial differential in total degree $\leq 10$.
We summarise the discussion in the following form, which will be sufficient for our later arguments.
Lemma 4.1.
Let $u_{5}\in E_{2}^{0,5}(k)\cong M^{5}$ be a nonzero class representing the fibre cohomology class of least positive degree. In total degree $j\leq 10$ all differentials $d_{r}(u_{5})$ vanish for $2\leq r\leq 5$ and for $r\geq 7$, and the only page on which $u_{5}$ can support a nontrivial differential is $r=6$. More precisely,
$d_{6}(u_{5})\in E_{6}^{6,0}(k)\cong H^{6}(BG_{2}).$
Fix once and for all a generator $x_{6}\in H^{6}(BG_{2})\cong\mathbb{F}_{2}$. Since $H^{6}(BG_{2})$ is one-dimensional over $\mathbb{F}_{2}$, there exists a uniquely determined scalar $\varepsilon(k)\in\mathbb{F}_{2}$ such that
$d_{6}(u_{5})=\varepsilon(k)\,x_{6}.$
Extending this analysis slightly, we obtain a convenient description of the possible differentials on the $H^{*}(BG_{2})$–submodule generated by $u_{5}$ in low total degrees.
Lemma 4.2.
Let $U\subset E_{2}^{*,*}(k)$ denote the $H^{*}(BG_{2})$–submodule generated by $u_{5}\in E_{2}^{0,5}(k)$. For classes in $U$ of total degree at most $10$ (equivalently, for classes $x\cdot u_{5}\in E_{2}^{s,5}(k)$ with $s+5\leq 10$), the following hold.
  1. (a)
    All differentials $d_{r}$ with $2\leq r\leq 5$ vanish on $U$.
  2. (b)
    Every differential
    $d_{6}\colon E_{6}^{s,5}(k)\longrightarrow E_{6}^{s+6,0}(k)$
    with $s+5\leq 10$ satisfies
    $d_{6}(x\cdot u_{5})=x\cdot d_{6}(u_{5})\qquad\text{for }x\in H^{s}(BG_{2}).$
  3. (c)
    All differentials $d_{r}$ with $r\geq 7$ vanish on $U$.
Proof.
By definition,
$U=H^{*}(BG_{2})\cdot u_{5}\subseteq E_{2}^{*,*}(k),$
so every class in $U$ of total degree $\leq 10$ has the form
$x\cdot u_{5}\in E_{2}^{s,5}(k),\qquad x\in H^{s}(BG_{2}),$
with $s+5\leq 10$.
Since
$H^{*}(BG_{2})\cong\mathbb{F}_{2}[x_{4},x_{6},x_{7}],$
the only values of $s$ for which $H^{s}(BG_{2})\neq 0$ and $s+5\leq 10$ are $s=0$ and $s=4$. Thus, in the range under consideration, the only possible bidegrees in $U$ are $(0,5)$ and $(4,5)$.
Proof of (a). Let $x\cdot u_{5}\in U$ with $s+5\leq 10$, and let $2\leq r\leq 5$. Then
$d_{r}(x\cdot u_{5})\in E_{r}^{s+r,\,5-r+1}(k).$
For $2\leq r\leq 5$ we have
$5-r+1\in\{4,3,2,1\}.$
By Proposition 3.2,
$M^{j}=H^{j}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})=0\qquad\text{for }1\leq j\leq 4.$
Hence the fibre term in bidegree (s+r,5r+1)(s+r,5-r+1) vanishes, so
$E_{r}^{s+r,\,5-r+1}(k)=0.$
Therefore
$d_{r}(x\cdot u_{5})=0\qquad\text{for all }2\leq r\leq 5.$
This proves (a).
Proof of (b). Now let $r=6$. Then
$d_{6}\colon E_{6}^{s,5}(k)\longrightarrow E_{6}^{s+6,0}(k)\cong H^{s+6}(BG_{2}).$
Since the Serre spectral sequence is multiplicative, the differential satisfies the Leibniz rule:
$d_{6}(x\cdot u_{5})=d_{6}(x)\cdot u_{5}+x\cdot d_{6}(u_{5}).$
But $x$ lies in bidegree $(s,0)$, so any differential on $x$ would land in bidegree (s+6,5)(s+6,-5), which is zero for degree reasons. Thus
$d_{6}(x)=0,$
and consequently
$d_{6}(x\cdot u_{5})=x\cdot d_{6}(u_{5}),$
as claimed. This proves (b).
Proof of (c). Let $r\geq 7$. For any class $x\cdot u_{5}\in U$ with $s+5\leq 10$, one has
$d_{r}(x\cdot u_{5})\in E_{r}^{s+r,\,5-r+1}(k).$
Since $r\geq 7$,
$5-r+1=6-r\leq-1.$
Thus the target has negative fibre degree, so
$E_{r}^{s+r,\,5-r+1}(k)=0.$
Hence
$d_{r}(x\cdot u_{5})=0\qquad\text{for all }r\geq 7.$
This proves (c). ∎
Note 4.3.
The case $k=0$ admits a particularly simple description. The class $0\in\pi_{4}(BG_{2})$ corresponds to the trivial principal $G_{2}$–bundle $S^{4}\times G_{2}\to S^{4}$, and by Gottlieb’s theorem [7], we have
$B\mathcal{G}_{0}\;\simeq\;\mathrm{Map}_{0}(S^{4},BG_{2}),$
where $\mathrm{Map}_{0}(S^{4},BG_{2})$ denotes the path component containing the constant map. The evaluation map $\mathrm{ev}\colon\mathrm{Map}_{0}(S^{4},BG_{2})\to BG_{2}$ admits a section $s\colon BG_{2}\to\mathrm{Map}_{0}(S^{4},BG_{2})$ defined by sending $y\in BG_{2}$ to the constant map at $y$. The existence of this section implies that the homomorphism
$\mathrm{ev}^{*}\colon H^{*}(BG_{2})\longrightarrow H^{*}(B\mathcal{G}_{0})$
is injective. Indeed, one has
$s^{*}\circ\mathrm{ev}^{*}=\mathrm{id}_{H^{*}(BG_{2})},$
so $\mathrm{ev}^{*}$ is split injective.
In the Serre spectral sequence for the evaluation fibration with $k=0$, the class $x_{6}\in H^{6}(BG_{2})\cong E_{2}^{6,0}$ must therefore survive to $E_{\infty}$ and cannot lie in the image of any differential. In particular, $x_{6}$ cannot be hit by $d_{6}(u_{5})$, forcing $d_{6}(u_{5})=0$. Thus, in our notation, $\varepsilon(0)=0$.
Lemma 4.4 (An 88–periodicity statement for $\varepsilon(k)$).
Work 22–locally. If $k\equiv k^{\prime}\pmod{8}$, then
$\varepsilon(k)=\varepsilon(k^{\prime})\in\mathbb{F}_{2}.$
In particular, $\varepsilon(k)$ depends only on the residue class of $k$ modulo $8$.
Proof.
By [11, Lemma 2.1], the connecting map
$\partial_{k}\colon G_{2}\to\Omega^{3}_{0}G_{2}$
is the triple adjoint of the Samelson product $\langle k\cdot i_{3},1\rangle$, and in particular
$\partial_{k}\simeq k\cdot\partial_{1}.$
Kameko proved that the 22–primary component of $\langle i_{3},1\rangle$ has order $4$ [10, Proposition 1.3]. Equivalently, after localization at $2$ the map $\partial_{1}$ has order $8$, so
$\partial_{k}\simeq\partial_{k^{\prime}}\qquad\text{$2$--locally whenever }k\equiv k^{\prime}\pmod{8}$.
Now the evaluation fibration
$\Omega^{3}_{0}G_{2}\longrightarrow B\mathcal{G}_{k}\xrightarrow{\;\mathrm{ev}\;}BG_{2}$
is a delooping of the homotopy fibration
$\mathcal{G}_{k}\longrightarrow G_{2}\xrightarrow{\;\partial_{k}\;}\Omega^{3}_{0}G_{2}.$
Hence, if $\partial_{k}\simeq\partial_{k^{\prime}}$ 22–locally, then the two homotopy fibrations over $G_{2}$ have equivalent homotopy fibres, and after delooping we obtain a fibre homotopy equivalence between the corresponding evaluation fibrations over $BG_{2}$. In particular, there is an isomorphism between the associated mod $2$ Serre spectral sequences which is compatible with the standard identifications
$E_{2}^{s,t}(k)\cong H^{s}(BG_{2})\otimes H^{t}(\Omega^{3}_{0}G_{2}),\qquad E_{2}^{s,t}(k^{\prime})\cong H^{s}(BG_{2})\otimes H^{t}(\Omega^{3}_{0}G_{2}).$
Under this identification, the class
$u_{5}\in E_{2}^{0,5}\cong H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$
corresponds to the same generator of the fibre cohomology, and
$x_{6}\in E_{2}^{6,0}\cong H^{6}(BG_{2};\mathbb{F}_{2})$
corresponds to the same degree-$6$ generator on the base. Therefore the $d_{6}$–differentials agree. Since $H^{6}(BG_{2};\mathbb{F}_{2})\cong\mathbb{F}_{2}\{x_{6}\}$ is one-dimensional, the scalar in
$d_{6}(u_{5})=\varepsilon(k)\,x_{6}$
must be the same for $k$ and $k^{\prime}$. Thus
$\varepsilon(k)=\varepsilon(k^{\prime}),$
as required. ∎
Corollary 4.5.
Work 22–locally. If $8\mid k$, then $\varepsilon(k)=0$.
Proof.
If $8\mid k$ then $k\equiv 0\pmod{8}$, so Lemma 4.4 gives $\varepsilon(k)=\varepsilon(0)$. By Note 4.3 we have $\varepsilon(0)=0$, hence $\varepsilon(k)=0$. ∎
4.3. The $E_{\infty}$–page and completion of the proof
For each total degree $j\leq 10$, the Serre spectral sequence yields a finite filtration
$0=F^{j+1}H^{j}(B\mathcal{G}_{k})\subseteq F^{j}H^{j}(B\mathcal{G}_{k})\subseteq\cdots\subseteq F^{0}H^{j}(B\mathcal{G}_{k})=H^{j}(B\mathcal{G}_{k})$
whose associated graded object is
$\mathrm{gr}\,H^{j}(B\mathcal{G}_{k})\cong\bigoplus_{s+t=j}E_{\infty}^{s,t}(k).$
By Lemmas 4.1 and 4.2, the contribution of the $H^{*}(BG_{2})$–submodule
$U=H^{*}(BG_{2})\cdot u_{5}\subseteq E_{2}^{*,*}(k)$
to $E_{\infty}^{*,*}(k)$ in total degree $j\leq 10$ is completely determined by the single differential
$d_{6}(u_{5})=\varepsilon(k)x_{6}.$
More precisely, within $U$ all differentials $d_{r}$ for $2\leq r\leq 5$ vanish, every $d_{6}$ is obtained from $d_{6}(u_{5})$ by multiplication with classes from $H^{*}(BG_{2})$, and no $d_{r}$ with $r\geq 7$ contributes in total degree $\leq 10$.
We do not claim here that these are the only possible differentials in total degree $\leq 10$ on the whole spectral sequence, since additional fibre classes in degrees $t\geq 6$ may in principle support further differentials. Thus Theorem 1.1 should be interpreted as a precise determination of the first kk–dependent contribution coming from the fibre class $u_{5}$.
Proof of Theorem 1.1.
Part (i) is the identification of the $E_{2}$–term with trivial local coefficients. Part (ii) is Proposition 3.2. Parts (iii) and (iv) follow from Lemmas 4.1 and 4.2. Part (v) is the standard filtration statement for a convergent cohomology Serre spectral sequence, together with the preceding description of the contribution of the submodule $U=H^{*}(BG_{2})\cdot u_{5}$. This completes the proof. ∎
Remark 4.6.
The restriction to total degree $j\leq 10$ in Theorem 1.1 is motivated by two factors. First, this range is sufficient to capture the first nontrivial cohomological dependence on the bundle class $k$, which manifests via the differential $d_{6}(u_{5})$. Second, extending the analysis beyond degree $10$ requires a substantially more detailed understanding of the fibre cohomology
$M=H^{*}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$
as an unstable $\mathcal{A}$–module, which would involve more elaborate Eilenberg–Moore spectral sequence computations beyond the scope of the present paper.
Once $M$ is determined in a larger range, the computation of $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ should reduce to a finite problem in the “hit” theory over the polynomial algebra $H^{*}(BG_{2})$, and may be amenable to computer implementation. It would also be interesting to compare the present low-degree analysis with Choi’s calculation [5] of the mod $2$ homology of the gauge groups $\mathcal{G}_{k}$ for $G=G_{2}$, via the path-loop fibration of $B\mathcal{G}_{k}$ and the resulting information on $H_{*}(\mathcal{G}_{k};\mathbb{F}_{2})$. We hope to return to these questions in future work.